OpenStudy (anonymous):

An equation in the system is y = x + 1. What is the other equation in the system if the solution is only at (0, 1)?

OpenStudy (anonymous):

A. x − y = 1 B. 2x + 2y = 4 C. x + y = 1 D. x + 2y = 2

OpenStudy (anonymous):

i thinl its A but unsure

OpenStudy (anonymous):

C. x + y = 1

OpenStudy (anonymous):

and also D. x + 2y = 2

OpenStudy (anonymous):

C